# ACTIVE: Continuous Similarity Search for Vessel Trajectories

### v2

@@ B.1.3. Effect of $l_{\textit{max}}$ and $l_{\text{min}}$
Figure 14. Effect of $l_{\text{min}}$ ($l_{\textit{max}}=l_{\text{min}}+20$)
@@
Figure 15. Effect of $l_{\textit{max}}$ ($l_{\text{min}}=30$)
@@
We also vary the maximum segment length $l_{\textit{max}}$, while fixing the minimum segment length $l_{\text{min}}$ at 30. The results are shown in Figure 15. An increase in $l_{\textit{max}}$ results in longer segmentation times for both datasets, as the process requires more iterations with larger $l_{\textit{max}}$ values. Additionally, the number of generated segments decreases slightly as $l_{\textit{max}}$ grows, leading to a modest reduction in both index size and the construction time during the building phase, i.e., DK-AIS-B and US-AIS-B.
B.2. Ablation Study
By default, CSTS incorporates segment pruning (Strategy 1), $k$-bound pruning (Strategy 2), and the incremental rule (Strategy 3). These three strategies are solely designed to improve efficiency and do not reduce the hit rate. In Figure 16, we report their impact on query time. Among the strategies, the incremental rule (S3) proves to be the most effective, reducing the query time by at least threefold. This is the evidence of the efficiency of the incremental rule in optimizing CSTS. The second-most effective strategy is $k$-bound pruning (S2), which shows comparable performance to S3. Segment pruning (S1), however, provides relatively less improvement on both datasets. This is because it introduces additional computational costs to compute the distance between a point and the minimum bounding rectangle of a segment.
The granularity adjustment (Strategy 4) is not applied in CSTS by default, as it may introduce errors. To evaluate the effectiveness of this strategy, we vary the granularity parameter $g$. Referring to Figure 17, both query time and hit rate decrease as $g$ increases. With a larger $g$, more points are disregarded during similarity computation, resulting in reduced query time. However, the hit rate declines because ignoring points introduces errors.
B.3. Case Study on Similarity Measures
Figure 18 presents case studies conducted on DK-AIS and US-AIS, illustrating the differences in results when applying different similarity measures. The orange line represents a moving object $o$, including both its current trajectory and the projected destination. The black line indicates the most similar trajectory identified using the OTRD measure. In comparison with other similarity measures such as Hausdorff, Fréchet, TrajCL, and CLEAR, the OTRD measure demonstrates a clear advantage. The other four measures focus predominantly on aligning with $o$’s current trajectory, resulting in matches that are less aligned with $o$’s overall path. Conversely, the trajectory identified by OTRD not only aligns closely with $o$’s current trajectory but also incorporates the future trend and destination, offering a more holistic similarity result. This highlights the strength of OTRD at capturing the dynamic and predictive aspects of trajectory comparison, making it suitable for applications where future trends and endpoints are important.
Figure 16. Effect of speed-up strategies
(a) Time vs. $g$
(b) Hit vs. $g$
Figure 17. Effect of granularity
Figure 18. Case study

### v3

@@ A-A4 Effect of $l_{\textit{max}}$ and $l_{\text{min}}$
Figure 17: Effect of $l_{\text{min}}$ ($l_{\textit{max}}=l_{\text{min}}+20$).
@@
Figure 18: Effect of $l_{\textit{max}}$ ($l_{\text{min}}=30$).
@@
We also vary the maximum segment length $l_{\textit{max}}$, while fixing the minimum segment length $l_{\text{min}}$ at 30. The results are shown in Fig. 18. An increase in $l_{\textit{max}}$ results in longer segmentation times for both datasets, as the process requires more iterations with larger $l_{\textit{max}}$ values. Additionally, the number of generated segments decreases slightly as $l_{\textit{max}}$ grows, leading to a modest reduction in both index size and the construction time during the building phase, i.e., DK-AIS-B and US-AIS-B.